Suppose that there were markets for two related goods. For Good 1, Demand is given by
Q1=198 -3P1 – 3P2 and Supply is given by Q1 = 4P1 -15. For Good 2, Demand is given by
Q2 = 294 – 4P2 – 2P1 and Supply given by Q2 = 4P2 – 20. From the information given,
determine the equilibrium price for Good 1 and Good 2.
1
Expert's answer
2020-12-21T03:38:22-0500

Solution:

The equilibrium price for Good 1:

D = S

"198-3P1-3P2=4P1-15"

"198+15-3P2=4P1+3P1"

"213-3P2=7P1"


"\frac{213}{3}-\frac{3P2}{3}=\frac{7P1}{3}"


"71-P2=2.33P1"


"P2=2.33P1-71"


The Equilibrium price for Good 2:

D = S

"294-4P2-2P1=4P2-20"

"294+20-2P1=4P2+4P2"

"314-2P1=8P2"


"\frac{314}{8}-\frac{2P1}{8}=\frac{8P2}{8}"


"39.25-0.25P1=P2"


"P2 = 0.25P1-39.25"